# A Longitudinal Study of Child Wellbeing Assessment via Online Interactions with a Social Robot

### v2

@@ 4.1.6 Measures
Post-session questionnaires: The user perception questionnaires were conducted by the experimenter after every experiment session. Total scores were computed for the first user perception questionnaire (Cronbach’s $\alpha$ = 0.65, mean score = 34.42 $\pm$ 3.74), inspired by [78] to understand children’s perception of the robot as a confidante, using the questionnaire ratings of 1: Strongly dislike, 2: Somewhat dislike, 3: Neither like nor dislike, 4: Somewhat like, and 5: Strongly like. For the second user perception questionnaire (Cronbach’s $\alpha$ = 0.84, mean score = 18.95 $\pm$ 8.27), inspired by the RAS [79] to understand the anxiety attributed to the robot by the children, total scores were computed using the response ratings of 1: I do not feel anxiety at all, 2: I hardly feel any anxiety, 3: I don’t feel much anxiety, 4: I feel a little anxiety, 5: I feel much anxiety, and 6: I feel anxiety very strongly. Finally, total scores from the third user perception questionnaire (Cronbach’s $\alpha$ = 0.82, mean score = 47.59 $\pm$ 3.33), inspired by the Godspeed questionnaire [80] to understand the user’s likeability of the robot and the perceived intelligence of the robot, were computed using the questionnaire ratings of 1- negative characteristic like "Unkind" to 5- positive characteristic like "Kind".
4.2 Construct Validity Analyses
In [81], we conducted validity analyses for our in-lab study described in [4]. We have followed the same approach for the data collected from the online remote study described in this work. We conducted construct validity analysis through confirmatory principal component analysis (PCA) to validate established psychological assessment tools such as the SMFQ and RCADS in longitudinal CRI research presented here. This was done to ensure they accurately measure the intended constructs and to comprehend their limitations within CRI contexts. Similarly, we applied the same analysis to both self-reported and parent-reported RCADS to achieve a deeper understanding of model fit and to discern the differences more clearly.
@@ 11. You thought nobody really loved you -
Table 1: Loadings of SMFQ Items
4.2.1 SMFQ
@@ 19. You/I/ My child have(has) problems with your/my/his or her appetite - - -
Table 2: Component Loadings for Robot-Administered, Self-Reported and Parent-Reported RCADS. The item wordings were adapted to the respective mode of administration
4.2.2 RCADS - Robot administered
@@
A confirmatory PCA was conducted with the 25 robot-administered RCADS items to examine the underlying structure. After inspecting the correlation matrix, it was confirmed that there was no threat of multicollinearity and that discriminant validity of variables was maintained, with the highest correlation of .59. The KMO measure of .831 affirmed sampling adequacy for the analysis, with all individual KMO values above .5. Bartlett’s test of sphericity, $\chi^{2}$ (300) = 1447.96, p $<$ .001, indicated the data was suitable for PCA. The PCA revealed one component with an eigenvalue of 8.79, explaining 35.16% of the variance. The scree plot justified retaining this single component. Table 2 shows the loadings of the different items of the robot-administered RCADS. High loadings were observed for several items, notably item 6 (.73) indicating strong relationships with the underlying factor, while the items with the lowest loadings was 6 (.46). Item 19 was not loaded, suggesting its loading was suppressed since it was below the .4 coefficient threshold considered significant for this analysis.
4.2.3 RCADS - Self reported
@@
A confirmatory PCA was conducted with the 25 self-reported RCADS items to examine the underlying structure. After inspecting the correlation matrix, it was confirmed that there was no threat of multicollinearity and that discriminant validity of variables was maintained, with the highest correlation of .83. The KMO measure of .65 affirmed sampling adequacy for the analysis, with all individual KMO values above .5. Bartlett’s test of sphericity, $\chi^{2}$ (300) = 682.01, p $<$ .001, indicated the data was suitable for PCA. The PCA revealed one component with an eigenvalue of 9.05, explaining 36.21% of the variance. The scree plot justified retaining this single component. Table 2 shows the loadings of the different items of the self-reported RCADS. High loadings were observed for several items, notably item 25 (.825) indicating strong relationships with the underlying factor, while the items with the lowest loading was item 14 (.45). Items 1, 5, 8, 15, and 19 were not loaded, suggesting their loadings were suppressed since they were below the .4 coefficient threshold considered significant for this analysis.
4.2.4 RCADS - Parents reported
@@
A confirmatory PCA was conducted with the 25 RCADS items to examine the underlying structure. After inspecting the correlation matrix, it was confirmed that there was no threat of multicollinearity and that discriminant validity of variables was maintained, with the highest correlation of .83. The KMO measure of .510 affirmed sampling adequacy for the analysis, with all individual KMO values above .5. Bartlett’s test of sphericity, $\chi^{2}$ (300) = 811.35, p $<$ .001, indicated the data was suitable for PCA. The PCA revealed one component with an eigenvalue of 8.04, explaining 32.17% of the variance. The scree plot justified retaining this single component. Table 2 shows the loadings of the different items of the parent-reported RCADS. High loadings were observed for several items, notably item 23 (.85) indicating strong relationships with the underlying factor, while the item with the lowest loadings was item 15 (.41). Items 8, 10, 11, 12, 19, and 25 were not loaded, suggesting their loadings were suppressed since they were below the .4 coefficient threshold considered significant for this analysis.
4.2.5 Transferbility of the measurements
@@
The construct validity analysis revealed a dominant principal component underlying each of the scales, suggesting that each scale includes a consistent factor for measuring children’s mental health in CRIs. However, not all items contributed equally to this factor. A comparison between robot-administered and self-reported RCADS highlighted the former’s superior construct validity (i.e., demonstrating a better fit), suggesting that mediating scale items via robotic interaction enhances the validity of symptom disclosure. This is in line with previous transferability analysis conducted with data from CRIs of similar structure [81]. The validity analysis not only elucidated the core factors measured by the scales in CRIs but also set the stage for the main analysis by establishing the superior construct validity of robot-administered assessments. This foundation enables a deeper exploration of how robotic mediation can improve psychological assessments. These results underpin the rationale for comparing wellbeing changes between tertiles for each session and across all sessions. The identification of a primary factor confirms that the scales can effectively distinguish between different levels of wellbeing measured using RCADS, making the Kruskal Wallis tests appropriate for identifying statistically significant differences between groups defined by tertiles of SMFQ scores. the validity analysis ensures that the instruments used in the study are measuring what they are intended to measure with a high degree of accuracy, thereby lending credibility and support to the subsequent formal analyses of between-subject and within-subject differences.
4.3 Data analyses
4.3.1 Data clustering.
@@ 4.3 Data analyses
We used the responses from the SMFQ to cluster the participants into three tertiles (low, med and high) similar to the framework followed in [4]. SMFQ thresholds were also taken into consideration in determining the children’s categorisation into the wellbeing clusters. For example, if tertile categorisation resulted in a child with an SMFQ score of 9 being placed in med tertile, we included the child in the high tertile as the threshold for SMFQ is a total score of 8  [^7] https://www.seattlechildrens.org/globalassets/documents/healthcare-professionals/pal/ratings/smfq-rating-scale.pdf. Tertile edges were computed separately for girls and boys in order to ensure similar participant numbers across the three wellbeing clusters (low tertile, med tertile and high tertile). Since the SMFQ is used to monitor symptoms of depression in children, children belonging to the low tertile and med tertile were very unlikely to obtain a diagnosis of depression while children in the high tertile were highly likely to obtain a diagnosis of depression.
4.3.2 Statistical analyses.
@@
Statistical analyses were performed using either the statistical toolbox in Matlab  [^8] https://uk.mathworks.com/products/statistics.html or using IBM SPSS  [^9] https://www.ibm.com/spss. Our primary analyses consisted of investigating wellbeing and user perception changes for the overall population (40 participants) across the varying levels of wellbeing (tertiles) as well as across the sessions. We also conducted a secondary exploratory analysis by further stratifying the data according to the gender of the participants. For this purpose, we employed non-parametric statistical analyses as our study data (RCADS, SMFQ and user perception responses) did not follow a normal distribution (Shapiro-Wilk test). We conducted Kruskal Wallis tests for between-subject analyses while investigating wellbeing or user perception changes between the tertiles for each session as well as across the tertiles for all the sessions combined. We employed Friedman’s test to conduct within-subject analyses while comparing changes in wellbeing and user perception responses longitudinally across the experiment sessions. Between-subject pairwise comparisons while investigating change in responses across girls and boys, between the children’s RCADS responses to the robot and the parent’s RCADS responses to the pre-study questionnaire, and across the remote study data collected in this work and our in-lab study [4, 62] were performed using Wilcoxon rank sum tests. Within-subject pairwise comparison between the children’s responses to the children’s RCADS responses to the robot and their self-report responses to the pre-study questionnaire was performed using the Wilcoxon Signed Rank test. Post-hoc tests for between-subject (Kruskal-Wallis test) and within-subject (Friedman’s test) analyses and the pairwise comparisons (across each investigated feature) were corrected using Bonferroni corrections ($0.05/3$). Incorrect responses were replaced by the closest response rating by the researchers (for ex. a "no" response to the SMFQ statements was replaced with "not true"). Missed responses to the statements of the questionnaires during the experiment session were imputed with the mode of the responses for the participant under consideration, for that questionnaire. In the case of three participants, data (two participants for the self-report questionnaire and one participant for the user perception questionnaires for session 3) did not get recorded due to technical issues and was replaced with mean value across all participants  [^10] https://www.theanalysisfactor.com/seven-ways-to-make-up-data-common-methods-to-imputing-missing-data/. However, the inputted data was not used to compute the reliability (Cronbach’s $\alpha$) and the construct validity of the questionnaires mentioned in Sec. 4.1.6 Measures and Sec. 4.2 Cunstruct Validity Analyses.
@@
Low tertile SMFQ <=3 15 SMFQ <=3 9 SMFQ <=4 8
Med tertile 3<SMFQ <=7 11 3<SMFQ <=7 5 4<SMFQ <=7 4
High tertile SMFQ >7 14 SMFQ > 7 6 SMFQ > 7 8
(b) In-lab study [4]
@@
Low tertile SMFQ <=2 16 SMFQ <=2 7 SMFQ <=2 9
Med tertile 2<SMFQ <=4 12 2 < SMFQ <=5 9 2 < SMFQ <= 5 6
High tertile SMFQ >4 13 SMFQ score >5 5 SMFQ >5 5
Table 3: Tertile edges and participant number (N) for the SMFQ scores for (a) remote study, and (b) In lab study [4].
5 Results
5.1 Tertile categorisation for identification of wellbeing clusters.
The SMFQ data collected during the first session was used for tertile categorisation of the participants into three wellbeing clusters similar to [4, 82, 83]. Table 3 shows the tertile edges and number of participants for the different population groups (overall, girls and boys) for both the remote study conducted in this work and our in-lab study [4], across the varying wellbeing clusters (low tertile, med tertile and high tertile).
Figure 3: (a) Longitudinal changes in wellbeing (RCADS scores) across different population groups (overall, girls and boys) categorised into wellbeing tertiles (low, med and high) using the SMFQ responses. (b) Overall changes in wellbeing (RCADS scores) across tertiles for all sessions combined. $*p<0.05$ corrected.
5.2 Effectiveness of online interactions with a social robot in wellbeing assessment (RQ1).
@@ 5 Results
Figure 4: RCADS responses across session 1 for the different modes of test administration from all population groups categorised into wellbeing tertiles (low, med and high) using the SMFQ responses. $*p<0.05$ corrected.
Since self-report responses and parent-report responses for the RCADS were collected before the experiment sessions, we have compared the responses of the RCADS across the three modes of test administration (self-report, parent-report and robot-report for session 1). As seen from Figure 4, Wilcoxon signed-rank test with a Bonferroni correction applied, has indicated a statistically significant difference between the self-report responses and the responses during session 1 with the robot for the high tertile ($Mdn_{self}=9,Mdn_{robot}=26,p=0.04$). Further, Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference between the parent-report responses and the responses during session 1 with the robot for the high tertile ($Mdn_{parent}=8.5,Mdn_{robot}=26,p=0.04$). We have not found any statistically significant differences across the tertiles for the RCADS responses between session 1 for the remote study with the robot and the RCADS responses for our in-lab study [4] for the overall population.
Figure 5: (a) Longitudinal changes in children’s anxiety measurement attributed to the robot across all population groups (overall, girls and boys) categorised into wellbeing tertiles (low, med and high) using the SMFQ responses. (b) Gender-based variability in the role of the robot as a confidante in the remote study. (c) Total percentage of the population that had consistent or improved user perception. User perception changes across tertiles for all sessions combined for (d) perception of the robot as a confidante and (e) robot anxiety measurement. $*p<0.05$ corrected.
5.3 The effect of the longitudinal interactions and co-location on user’s perception of the robot (RQ2).
As seen in Figure 5, Friedman’s test indicated a statistically significant difference in the robot anxiety measurement scores for the overall population group belonging to the low tertile, across the three sessions ($\chi^{2}(2)=6.87,p=0.032$). Post hoc analysis with Wilcoxon signed-rank tests was conducted with a Bonferroni correction applied, resulting in statistical significance between the robot anxiety measurement responses for session 1 and session 3 ($Mdn_{session1}=16,Mdn_{session2}=14,p=0.041$). We have observed (Figure 5 (c)) that out of 40 participants (20 girls and 20 boys), most participants have reported an average increase or have sustained their opinion of the robot’s likeability, perceived intelligence and role as a confidante. In the case of robot anxiety, most participants have reported an average decrease in robot anxiety or have maintained their opinion of the robot across the three sessions.
We did not find any statistically significant difference in the responses to the perception questionnaires (robot as confidante, robot anxiety measurement and likeability and perceived intelligence of the robot) across the tertiles throughout any of the sessions (session 1, session 2, and session 3). However, we have observed (Figure 5 (d)) that the Kruskal Wallis H test has shown a statistically significant difference ($H(2)=11.13,p=0.004$) in the children’s perception of the robot as a confidante across tertiles for all the sessions combined. Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=33,Mdn_{high}=37,p=0.004$). Kruskal Wallis H test has also shown a statistically significant difference ($H(2)=8.82,p=0.012$) in the robot anxiety measurement across tertiles for all the sessions combined (Figure 5 (e)). Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=15,Mdn_{high}=21,p=0.012$). We did not find any statistical significance between tertiles for all sessions combined in the case of children’s likeability of the robot and their perceived intelligence of the robot.
As seen from Figure 6 (a), Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference for the robot anxiety measurement in session 1 of the remote study and our in-lab study [62] for the med tertile of the overall population group ($Mdn_{remote}=18,Mdn_{in-lab}=24,p=0.038$). Further, Wilcoxon rank sum test (Figure 6 (b)) with a Bonferroni correction applied has also indicated a statistically significant difference for the perceived intelligence of the robot in session 1 of the remote study and our in-lab study [62] for the medium tertile of the overall population group ($Mdn_{remote}=24,Mdn_{in-lab}=21,p=0.045$). Wilcoxon rank sum test with a Bonferroni correction applied has also indicated a statistically significant difference between the perceived intelligence of the robot in session 1 of the remote study and our in-lab study [62] for the high tertile of the overall population group ($Mdn_{remote}=24,Mdn_{in-lab}=20.5,p=0.045$).
Figure 6: User perception changes across the remote study and our in-lab study [62] categorised into wellbeing tertiles (low, med and high) using the SMFQ responses. $*p<0.05$ corrected.
5.4 Gender-based differences in the effectiveness of longitudinal wellbeing assessment and user perception (RQ3).
@@
From our exploratory supplementary analyses, we have observed that in the case of girls (Figure 3 (a)), Kruskal Wallis H test has shown a statistically significant difference ($H(2)=10.64,p=0.005$) in the RCADS responses for session 1 between the tertiles. Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=6,Mdn_{high}=28,p=0.006$) and between med tertile and high tertile ($Mdn_{med}=6,Mdn_{high}=28,p=0.042$). Further, Friedman’s test indicated a statistically significant difference (Figure 3(a)) in the RCADS scores for girls belonging to the high tertile, across the three sessions ($\chi^{2}(2)=11.14,p=0.004$). Post hoc analysis with Wilcoxon signed-rank tests was conducted with a Bonferroni correction applied, resulting in statistical significance between the RCADS responses for session 1 and session 2 ($Mdn_{session1}=28,Mdn_{session2}=11.5,p=0.007$).
@@
In the case of boys, as seen in Figure 3 (a), Kruskal Wallis H test has shown a statistically significant difference ($H(2)=7.91,p=0.019$) in the RCADS responses for session 1 between the tertiles. Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=9,Mdn_{high}=23,p=0.016$). Kruskal Wallis H test has also shown a statistically significant difference ($H(2)=9.22,p=0.010$) in the RCADS responses for session 2 between the tertiles. Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=8.5,Mdn_{high}=20.5,p=0.009$). We did not find any statistically significant difference longitudinally (between session 1, session 2, and session 3) between the RCADS responses for boys across all tertiles. We have also not observed any statistically significant differences (Figure 4) between the modes of RCADS administration (robot report, self-report, and parent-report) across all the tertiles for girls. In the case of boys (Figure 4) Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference between the parent-report responses and the responses during session 1 with the robot for the high tertile ($Mdn_{parent}=7,Mdn_{robot}=23,p=0.013$). Further, even in the case of girls and boys, we have not found any statistically significant differences across the tertiles for the RCADS responses between session 1 for the remote study with the robot and the RCADS responses for our in-lab study presented in [4].
We also did not find any statistical significance while comparing user perception responses to the questionnaires for girls or boys, across the tertiles, throughout any of the sessions. However, Wilcoxon rank sum test (Figure 5) has indicated a statistically significant difference between the children’s perception of the robot as a confidante for the high tertile between girls and boys ($Mdn_{girls}=37,Mdn_{boys}=33.5,W=67,p=0.012$). While considering user perception changes for girls (Figure 6(c)), Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference for the perception of the robot as a confidante in session 1 of the remote study and our in-lab study [62] for the medium tertile of the overall population group ($Mdn_{remote}=37,Mdn_{in-lab}=31,p=0.038$). Further, in the case of boys (Figure 6(d)), Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference for robot anxiety measurement in session 1 of the remote study and our in-lab study [62] for the med tertile of the overall population group ($Mdn_{remote}=20,Mdn_{in-lab}=24,p=0.048$).

### v3

@@ 3.2.6. Measures
Post-session questionnaires: The user perception questionnaires were conducted by the experimenter after every experiment session. Total scores were computed for the first user perception questionnaire (Cronbach’s $\alpha$ = 0.65, mean score = 34.42 $\pm$ 3.74), inspired by (Murray et al., 2022) to understand children’s perception of the robot as a confidante, using the questionnaire ratings of 1: Strongly dislike, 2: Somewhat dislike, 3: Neither like nor dislike, 4: Somewhat like, and 5: Strongly like. For the second user perception questionnaire (Cronbach’s $\alpha$ = 0.84, mean score = 18.95 $\pm$ 8.27), inspired by the RAS (Nomura et al., 2006) to understand the anxiety attributed to the robot by the children, total scores (referred to as ”robot anxiety”) were computed using the response ratings of 1: I do not feel anxiety at all, 2: I hardly feel any anxiety, 3: I don’t feel much anxiety, 4: I feel a little anxiety, 5: I feel much anxiety, and 6: I feel anxiety very strongly. Finally, total scores from the third user perception questionnaire (Cronbach’s $\alpha$ = 0.82, mean score = 47.59 $\pm$ 3.33), inspired by the Godspeed questionnaire (Bartneck et al., 2009) to understand the user’s likeability of the robot and the perceived intelligence of the robot, were computed using the questionnaire ratings of 1- negative characteristic like ”Unkind” to 5- positive characteristic like ”Kind”.
3.3. Construct Validity Analyses
In (Abbasi et al., 2024a), we conducted validity analyses for our in-lab study described in (Abbasi et al., 2022, 2024b). We have followed the same approach for the data collected from the online remote study described in this work. We conducted construct validity analysis through confirmatory principal component analysis (PCA) to validate established psychological assessment tools such as the SMFQ and RCADS in longitudinal CRI research presented here. This was done to ensure they accurately measure the intended constructs and to comprehend their limitations within CRI contexts. Similarly, we applied the same analysis to both self-reported and parent-reported RCADS to achieve a deeper understanding of model fit and to discern the differences more clearly.
@@ 11. You thought nobody really loved you -
Table 1. Loadings of SMFQ Items
3.3.1. SMFQ
@@ 19. You/I/ My child have(has) problems with your/my/his or her appetite - - -
Table 2. Component Loadings for Robot-Administered, Self-Reported and Parent-Reported RCADS. The item wordings were adapted to the respective mode of administration
3.3.2. RCADS - Robot administered
@@
A confirmatory PCA was conducted with the 25 robot-administered RCADS items to examine the underlying structure. After inspecting the correlation matrix, it was confirmed that there was no threat of multicollinearity and that discriminant validity of variables was maintained, with the highest correlation of .59. The KMO measure of .831 affirmed sampling adequacy for the analysis, with all individual KMO values above .5. Bartlett’s test of sphericity, $\chi^{2}$ (300) = 1447.96, p $<$ .001, indicated the data was suitable for PCA. The PCA revealed one component with an eigenvalue of 8.79, explaining 35.16% of the variance. The scree plot justified retaining this single component. Table 2 shows the loadings of the different items of the robot-administered RCADS. High loadings were observed for several items, notably item 6 (.73) indicating strong relationships with the underlying factor, while the items with the lowest loadings was 25 (.45). Item 19 was not loaded, suggesting its loading was suppressed since it was below the .4 coefficient threshold for this analysis.
3.3.3. RCADS - Self reported
@@
A confirmatory PCA was conducted with the 25 self-reported RCADS items to examine the underlying structure. After inspecting the correlation matrix, it was confirmed that there was no threat of multicollinearity and that discriminant validity of variables was maintained, with the highest correlation of .83. The KMO measure of .65 affirmed sampling adequacy for the analysis, with all individual KMO values above .5. Bartlett’s test of sphericity, $\chi^{2}$ (300) = 682.01, p $<$ .001, indicated the data was suitable for PCA. The PCA revealed one component with an eigenvalue of 9.05, explaining 36.21% of the variance. The scree plot justified retaining this single component. Table 2 shows the loadings of the different items of the self-reported RCADS. High loadings were observed for several items, notably item 25 (.83) indicating strong relationships with the underlying factor, while the items with the lowest loading was item 14 (.45). Items 1, 5, 8, 15, and 19 were not loaded, suggesting their loadings were suppressed since they were below the .4 coefficient threshold for this analysis.
3.3.4. RCADS - Parents reported
@@
A confirmatory PCA was conducted with the 25 RCADS items to examine the underlying structure. After inspecting the correlation matrix, it was confirmed that there was no threat of multicollinearity and that discriminant validity of variables was maintained, with the highest correlation of .83. The KMO measure of .510 affirmed sampling adequacy for the analysis, with all individual KMO values above .5. Bartlett’s test of sphericity, $\chi^{2}$ (300) = 811.35, p $<$ .001, indicated the data was suitable for PCA. The PCA revealed one component with an eigenvalue of 8.04, explaining 32.17% of the variance. The scree plot justified retaining this single component. Table 2 shows the loadings of the different items of the parent-reported RCADS. High loadings were observed for several items, notably item 23 (.85) indicating strong relationships with the underlying factor, while the item with the lowest loadings was item 15 (.41). Items 8, 10, 11, 12, 19, and 25 were not loaded, suggesting their loadings were suppressed since they were below the .4 coefficient threshold for this analysis.
3.3.5. Transferbility of the measurements
@@
The construct validity analysis revealed a dominant principal component underlying each of the scales, suggesting that each scale includes a consistent factor for measuring children’s mental health in CRIs. However, not all items contributed equally to this factor. A comparison between robot-administered and self-reported RCADS highlighted the former’s superior construct validity (i.e., demonstrating a better fit), suggesting that mediating scale items via robotic interaction enhances the validity of symptom disclosure. This is in line with previous transferability analysis conducted with data from CRIs of similar structure (Abbasi et al., 2024a). The validity analysis not only elucidated the core factors measured by the scales in CRIs but also set the stage for the main analysis by establishing the superior construct validity of robot-administered assessments. This foundation enables a deeper exploration of how robotic mediation can improve psychological assessments. These results underpin the rationale for comparing wellbeing changes between tertiles for each session and across all sessions. The identification of a primary factor confirms that the scales can effectively distinguish between different levels of wellbeing measured using RCADS, making the Kruskal Wallis tests appropriate for identifying statistically significant differences between groups defined by tertiles of SMFQ scores. The construct validity analysis ensures that the instruments used in the study are measuring what they are intended to measure with a high degree of accuracy, thereby lending credibility and support to the subsequent formal analyses of between-subject and within-subject differences.
3.4. Data analyses
3.4.1. Data clustering.
@@ 3.4. Data analyses
We used the responses from the SMFQ to cluster the participants into three tertiles (low, med and high) similar to the framework followed in (Abbasi et al., 2022, 2024b). SMFQ thresholds were also taken into consideration in determining the children’s categorisation into the wellbeing clusters. For example, if tertile categorisation resulted in a child with an SMFQ score of 9 being placed in med tertile, we included the child in the high tertile as the threshold for SMFQ is a total score of 8  [^4] https://www.seattlechildrens.org/globalassets/documents/healthcare-professionals/pal/ratings/smfq-rating-scale.pdf. Tertile edges were computed separately for girls and boys in order to ensure similar participant numbers across the three wellbeing clusters (low tertile, med tertile and high tertile). Since the SMFQ is used to monitor symptoms of mood disorders in children, children belonging to the low tertile and med tertile were very unlikely to obtain a diagnosis of mood disorders while children in the high tertile were more likely to obtain a diagnosis of mood disorders.
Following the tertile categorisation, these well-being clusters were used as independent variables to analyse their relationship with outcomes assessed by the RCADS. This approach, well-supported in mental health research, facilitates structured comparisons of how varying levels of mood, as indicated by SMFQ tertiles, influence specific outcomes such as anxiety and depression symptoms (Shields et al., 2023; Johnson et al., 2021). Given the distinct yet related constructs measured by SMFQ (general mood) and RCADS (specific anxiety and depression symptoms), this method allows us to assess how mood categories align with different levels of psychological symptoms. While a correlation between SMFQ and RCADS is typically expected (e.g., Lisøy et al., 2022), the CRI context, where the robot administered RCADS, required careful consideration, as we could not assume this correlation would hold in the same way as in traditional child mental health studies.
3.4.2. Statistical analyses.
@@
Statistical analyses were performed using either the statistical toolbox in Matlab  [^5] https://uk.mathworks.com/products/statistics.html or using IBM SPSS  [^6] https://www.ibm.com/spss. Our primary analyses consisted of investigating wellbeing and user perception changes for the overall population (40 participants) across the varying levels of wellbeing (tertiles) as well as across the sessions. We also conducted a secondary exploratory analysis by further stratifying the data according to the gender of the participants. For this purpose, we employed non-parametric statistical analyses as our study data (RCADS, SMFQ and user perception responses) did not follow a normal distribution (Shapiro-Wilk test). We conducted Kruskal Wallis tests for between-subject analyses while investigating wellbeing or user perception changes between the tertiles for each session as well as across the tertiles for all the sessions combined. We employed Friedman’s test to conduct within-subject analyses while comparing changes in wellbeing and user perception responses longitudinally across the experiment sessions. Between-subject pairwise comparisons while investigating change in responses across girls and boys, between the children’s RCADS responses to the robot and the parent’s RCADS responses to the pre-study questionnaire, and across the remote study data collected in this work and our in-lab study (Abbasi et al., 2022, 2023, 2024b) were performed using Wilcoxon rank sum tests. Within-subject pairwise comparison between the children’s responses to the children’s RCADS responses to the robot and their self-report responses to the pre-study questionnaire was performed using the Wilcoxon Signed Rank test.
To control for Type I errors due to multiple hypothesis testing, we applied the Holm-Bonferroni correction to p-values from the confirmatory hypothesis tests. This method sequentially adjusts significance levels for each hypothesis, prioritising smaller p-values to rigorously control Type I error while preserving sensitivity (Aickin and Gensler, 2011). Both corrected and uncorrected p-values are reported in the results. Post-hoc tests for between-subject (Kruskal-Wallis test) and within-subject (Friedman’s test) analyses and the pairwise comparisons (across each investigated feature) were corrected using Bonferroni corrections ($0.05/3$).
Incorrect responses were replaced by the closest response rating by the researchers (e.g., a ”no” response to the SMFQ statements was replaced with ”not true”). Missed responses to the statements of the questionnaires during the experiment session were imputed with the mode of the responses for the participant under consideration, for that questionnaire. In the case of three participants, data (two participants for the self-report questionnaire and one participant for the user perception questionnaires for session 3) did not get recorded due to technical issues and was replaced with mean value across all participants [^7] https://www.theanalysisfactor.com/seven-ways-to-make-up-data-common-methods-to-imputing-missing-data/. However, the inputted data was not used to compute reliability scores (Cronbach’s $\alpha$) and to evaluate the construct validity of the measurements mentioned in Sec. 3.2.6 Measures and Sec. 3.3 Construct Validity Analyses.
@@
Low tertile SMFQ ¡=3 15 SMFQ ¡=3 9 SMFQ ¡=4 8
Med tertile 3¡SMFQ ¡=7 11 3¡SMFQ ¡=7 5 4¡SMFQ ¡=7 4
High tertile SMFQ ¿7 14 SMFQ ¿ 7 6 SMFQ ¿ 7 8
(b) In-lab study (Abbasi et al., 2022, 2024b)
@@
Low tertile SMFQ ¡=2 16 SMFQ ¡=2 7 SMFQ ¡=2 9
Med tertile 2¡SMFQ ¡=4 12 2 ¡ SMFQ ¡=5 9 2 ¡ SMFQ ¡= 5 6
High tertile SMFQ ¿4 13 SMFQ score ¿5 5 SMFQ ¿5 5
Table 3. Tertile edges and participant number (N) for the SMFQ scores for (a) remote study, and (b) In lab study (Abbasi et al., 2022, 2024b).
4. Results
4.1. Tertile categorisation for identification of wellbeing clusters.
The SMFQ data collected during the first session was used for tertile categorisation of the participants into three wellbeing clusters similar to (Abbasi et al., 2022; Jachens and Houdmont, 2019; Munir et al., 2014). Table 3 shows the tertile edges and number of participants for the different population groups (overall, girls and boys) for both the remote study conducted in this work and our in-lab study (Abbasi et al., 2022, 2024b), across the varying wellbeing clusters (low tertile, med tertile and high tertile).
Figure 3. (a) Longitudinal changes in wellbeing (RCADS scores: dependent variables) across different population groups (overall, girls and boys) categorised into wellbeing tertiles (low, med and high; independent variables) using the SMFQ responses. (b) Children’s wellbeing (RCADS scores) across tertiles for all sessions combined. $*p<0.05$ corrected.
4.2. Effectiveness of online interactions with a social robot in wellbeing assessment (RQ1).
@@ 4. Results
In the case of the RCADS responses for session 2, Kruskal Wallis H test has shown a statistically significant difference ($H(2)=10.73,p=0.005$; $p=0.010$, Holm-Bonferroni corrected) between the tertiles. Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=7,Mdn_{high}=18.5,p=0.004$).
@@
While considering the RCADS responses for session 3, Kruskal Wallis H test has also shown a statistically significant difference ($H(2)=12.02,p=0.002$; $p=0.010$, Holm-Bonferroni corrected) between the tertiles. Post-hoc pairwise comparisons using Dunn’s tests, with a Bonferroni correction applied, have indicated statistically significant differences between low tertile and high tertile ($Mdn_{low}=6,Mdn_{high}=18.5,p=0.004$).
[…]
Figure 4. RCADS responses (dependent variables) across session 1 for the different modes of test administration from all population groups categorised into wellbeing tertiles (low, med and high; independent variables) using the SMFQ responses. $*p<0.05$ corrected.
Since self-report responses and parent-report responses for the RCADS were collected before the experiment sessions, we have compared the responses of the RCADS across the three modes of test administration (self-report, parent-report and robot-report for session 1). As seen from Figure 4, Wilcoxon signed-rank test with a Bonferroni correction applied, has indicated a statistically significant difference between the self-report responses and the responses during session 1 with the robot for the high tertile ($Mdn_{self}=9,Mdn_{robot}=26,p=0.04$). Further, Wilcoxon rank sum test with a Bonferroni correction applied has indicated a statistically significant difference between the parent-report responses and the responses during session 1 with the robot for the high tertile ($Mdn_{parent}=8.5,Mdn_{robot}=26,p=0.04$). We have not found any statistically significant differences across the tertiles for the RCADS responses between session 1 for the remote study with the robot and the RCADS responses for our in-lab study (Abbasi et al., 2022, 2024b) for the overall population.